OpenStudy (anonymous):

What is the 5th partial sum of sum of -2+5n infinity

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} -2+5n\]

jhonyy9 (jhonyy9):

will be -2+5*1 +(-2)+5*2 +(-2)+5*3 ... so because there have wrote partial sum i think that will be -2+5*5=-2+25 =23

OpenStudy (anonymous):

Thank you